OpenStudy (anonymous):

Why is the domain and range of x^2+y^2=1 this answer?

OpenStudy (anonymous):

\[D = {-1\le1 \le1}\]

OpenStudy (anonymous):

maybe \[-1\leq x\leq 1\]

OpenStudy (turingtest):

you mean\[-1\le x\le 1\]what if x=2 ? what would y be?

OpenStudy (anonymous):

\[R = -1 \le y \le 1\]

OpenStudy (anonymous):

ohh I meant that

OpenStudy (anonymous):

yes you are right, it is the unit circle

OpenStudy (anonymous):

I get so confused with domains and ranges D:

OpenStudy (anonymous):

like why is it -1 less than or equal to x less than or equal to 1?

OpenStudy (turingtest):

what if \[|y|>1\]what would x be?

OpenStudy (anonymous):

i don't know.

OpenStudy (turingtest):

try it with x=2 what would y be?

OpenStudy (anonymous):

what do you mean?

OpenStudy (experimentx):

can't call it a function though ... it's a closed curve.

OpenStudy (anonymous):

I just don't understand why it's greater than or less than.. o.o

OpenStudy (turingtest):

@milliex51 can you solve the above for y ? an experimX is correct as usual, but that doesn't really matter for what we're talking about

OpenStudy (turingtest):

solve\[x^2+y^2=1\]for \(y\) please :)

OpenStudy (anonymous):

\[y = +/- \sqrt{-x}+1\]

OpenStudy (anonymous):

-x^2

OpenStudy (turingtest):

not quite...

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

or do I close bracket the 1 too?

OpenStudy (turingtest):

\[y=\pm\sqrt{1-x^2}\]

OpenStudy (anonymous):

ohh..

OpenStudy (turingtest):

\[x^2+y^2=1\]\[y^2=1-x^2\]\[y=\pm\sqrt{1-x^2}\]now what if x=2 ? what do you get for y ?

OpenStudy (experimentx):

now if we redefine it x^2+y^2=1 by y = sqrt(1-x^2) & y = -sqrt(1-x^2) on a field of real numbers than ..

OpenStudy (turingtest):

^echo, echo.... :P

OpenStudy (experimentx):

TuringTest was thinking exactly same as me

OpenStudy (anonymous):

isn't it the same from my answer? um, x=2 would be y=non real?

OpenStudy (turingtest):

and if y is not real, then x=2 is not in the domain, right?

OpenStudy (anonymous):

right.

OpenStudy (turingtest):

what about any \[|x|>1\]? what will y be?

OpenStudy (anonymous):

how do I solve that?

OpenStudy (turingtest):

I asking you if you plug in any number greater than 1 in for x into the formula we have for y, what will y e?

OpenStudy (turingtest):

*what will y be?

OpenStudy (anonymous):

i don't know D:

OpenStudy (turingtest):

ok, look at it this way...

OpenStudy (turingtest):

what is under the radical in\[y=\pm\sqrt{1-x^2}\]cannot be negative, right? otherwise y would be imaginary. do we agree?

OpenStudy (anonymous):

yes, yup.

OpenStudy (turingtest):

so that mean that we have a requirement that\[1-x^2\ge0\]right?

OpenStudy (anonymous):

yup. or else it will be undefined

OpenStudy (anonymous):

or imaginary

OpenStudy (turingtest):

so solve that requirement for x and you will get your domain

OpenStudy (anonymous):

but how? ahhh

OpenStudy (turingtest):

\[1-x^2\ge0\]\[1\ge x^2\]perhaps here is where you get a little confused...\[1\ge x\ge -1\]

OpenStudy (anonymous):

ohh.. I transpose the negative x^2

OpenStudy (turingtest):

remember that x^2 is always positive, so all that matters is that\[|x|\le1\]and y will be real

OpenStudy (anonymous):

how about when I have to look at a graph?

OpenStudy (anonymous):

Like:

OpenStudy (anonymous):

what signs am I going to use? o.o

OpenStudy (turingtest):

we are saying that x can take on any value in the black region, right?